OpenStudy (anonymous):

it takes Kevin 6 1/2 minutes to walk 8 1/4 blocks. At this rate, how many blocks can he walk in 1 minute???

OpenStudy (snowsurf):

Well we know that he walking at some rate. Lets call it r1. Then the problem ask if he is walking at that same rate how many blocks would he had moved. That rate will be r2. Since the problem says at this rate we know that: \[r_1=r_2\] And you know that d = rt where d is distance r is the rate and t is the time. If we solve for r and set both r equal to each other then we have: \[\frac{ d_1 }{ t_1}=\frac{ d_2}{ t_2 }\] Now substitute values in that are given \[\frac{ 8\frac{ 1 }{ 4 } }{ 6\frac{ 1 }{ 2 } }=\frac{ x }{ 1 }\] Just solve for x